Solve the system by addition.
5x – 3y = 13
4x – 3y = 11

Answer by jsteppergurl(2)

Okay. First, you're going to need one of the three terms to add up with the other term and make 0. Since none of them do this, you're going to have to multiply or divide them. -3y+3y=0, so you're going to want to multiply the bottom equation by negative 1. (-1) So (-1)4x-(-1)3y=(-1)11. Simplify this and you get -4x+3y=-11. Now you're going to add this system straight down, but use the new bottom equation. Also, every time you have a minus sign it is the same thing as plus a negative number. So 2-1 is the same thing as 2+(-1). So you have...
5x + (-3y) = 13
+
-4x + 3y = -11
--------------------
x = 2
just add straight down so 5x+(-4x)=x
(-3y)+3y=0
and 13+(-11)=2
x=2